OpenStudy (loser66):

Please, help Prove \(f(x) =\dfrac{x}{(\sqrt (x^2+1)}\) is onto from R to (-1,1)

OpenStudy (loser66):

The previous part is f(x) is one to one and I got it. Just this part I need to confirm

OpenStudy (ikram002p):

what u got so far ?

OpenStudy (loser66):

So far, I get : let y in (-1,1) , then \(\dfrac{x}{\sqrt{x^2+1}=y\)

OpenStudy (loser66):

if and only if \(x^2= y^2(x^2+1)\) --> \(x^2 -y^2x^2 -y^2 =0\)

OpenStudy (loser66):

--> \(x^2(1-y^2) -y^2=0\) so that \(x= \pm \dfrac{y}{\sqrt{1-y^2}}\)

OpenStudy (loser66):

and ... stuck

OpenStudy (amistre64):

can we determine a min and max by chance? not that its proof, but could be usefl

OpenStudy (amistre64):

take the limit as x to +- infinity

OpenStudy (amistre64):

we can observe that there is no real value that will mess up the function; ie 0 on the bottom, or negative sqrt

OpenStudy (amistre64):

if the slope if the function is always postive, and limits to 0 at the ends ...

OpenStudy (amistre64):

well, always postive or always negative ... monotonic is the term maybe?

OpenStudy (loser66):

so??

OpenStudy (rational):

For onto, you need to prove y takes all the values between (-1,1)

OpenStudy (amistre64):

\[\lim_{x\to -inf}\frac{x}{\sqrt{x^2+1}}=-1\] \[\lim_{x\to +inf}\frac{x}{\sqrt{x^2+1}}=1\] \[f'=\frac{1}{(x^2+1)^{3/2}}\] f' is never 0, and always positive

OpenStudy (rational):

derivative test looks neat than finding x and bounding y

OpenStudy (loser66):

@amistre64 that is the part I prove f is one to one

OpenStudy (amistre64):

if you go that route, try contradiction

OpenStudy (amistre64):

it proves onto to me, since there is not x in R that is restricted, and that y = -1, y=0, y=1 are withing the range limits

OpenStudy (rational):

that proves onto also because the lower and upper limits are proven to be -1 and +1

OpenStudy (rational):

two birds in one shot !

OpenStudy (amistre64):

f'(+- inf) to 0 as well

OpenStudy (amistre64):

might not be book correct, but i like it :)

OpenStudy (loser66):

Please show me how \[\lim_{x\to -inf}\frac{x}{\sqrt{x^2+1}}=-1\]

OpenStudy (rational):

\(\large x^2+1 \approx x^2\) as \(\large x\to \infty \)

OpenStudy (loser66):

\[\lim_{x\to -inf}\frac{x}{\sqrt{x^2+1}}=\lim_{x\to -inf}\frac{x}{x\sqrt{1+1/x^2}}=\lim_{x\to -inf}\frac{1}{\sqrt{1+0}}\] how to get -1

OpenStudy (rational):

change of variables y = -x

OpenStudy (rational):

\[\large \lim_{x\to -\infty}\frac{x}{\sqrt{x^2+1}}= \lim_{y\to \infty}\frac{-y}{\sqrt{y^2+1}}\]

OpenStudy (loser66):

Thank you.

OpenStudy (loser66):

Do we have to show f(x) is continuous on the interval before taking lim?

OpenStudy (ikram002p):

A function is onto when its image equals its range so you need to show that :- 1_ domain and range from (-1,1) 2_ both of f(x) and f^-1(x) are continues

OpenStudy (loser66):

One more question: to have f is onto from R to (-1,1) , we have to show that for all y in (-1,1) , x is defined in R but, when solving \( y = \dfrac{x}{\sqrt{x^2+1}}\) in the interval (-1,1) , it means \(-1<\dfrac{x}{\sqrt{x^2+1}}<1\) and I can't define x satisfy that expression

OpenStudy (loser66):

I see nothing, girl :)

OpenStudy (anonymous):

Maybe you could say $$\frac{x}{\sqrt{x^2+1}}=y$$and solve for x? This yields: $$x=\sqrt{\frac{y^2}{1-y^2}}$$ which is only a real number if: $$-1<y<1.$$

OpenStudy (loser66):

Best answer!! clear and neat.

OpenStudy (loser66):

Thank you so much. It makes sense to me.

OpenStudy (anonymous):

You're welcome :)